how many gallons of a 25% acid solution must be mixed with 100 gallons of a 6% acid solution to obtain a mixture that is 20% acid?
Answer by stanbon(75887)

how many gallons of a 25% acid solution must be mixed with 100 gallons of a 6% acid solution to obtain a mixture that is 20% acid?
------------------
Equations:
acid + acid = acid
---
0.25x + 0.06*100 = 0.20(x+100)
------
Multiply thru by 100 to get:
25x + 6*100 = 20x+20*100
5x = 14*100
x = 14*20
x = 280 gallons (amt of 25% solution needed)
===============================================
Cheers,
Stan H.